# Exploring the Impact of Dataset Statistical Effect Size on Model Performance and Data Sample Size Sufficiency

### v2

@@ Abstract
Codehttps://colab.research.google.com/drive/17cZ8pWYAQzCcENLgoUS01cRj5uLWi8ta?usp=sharing
Datasetshttp://archive.ics.uci.edu/dataset/2/adult
Introduction
@@
The sufficiency of a dataset, both in terms of size and representativeness, is critical to training an effective Machine Learning model. Failure to train on data of adequate quality is likely to result in models that dramatically under-perform in production environments (Friedland 2024).
@@
Training machine learning models on inadequate data presents several challenges that can significantly impact model performance and generalizability. While the problem of ’inadequacy’ of data is a multifaceted one, one significant component of it is a lack of sufficient data volume, which in-turn likely impacts the representativeness of the data sample towards the overall population undermining the ability of a model to learn meaningful patterns, resulting in overfitting, where the model performs well on the training data but struggles to generalize to unseen data leading to poor outcomes when deployed in real-world scenarios. This owes to the fact that if the dataset used for training does not accurately reflect the conditions and distributions the model will encounter in real-world applications, the model will struggle to generalize beyond the narrow scope of its training (Masiha et al. 2021). Consequently, the model’s assumptions about the domain may be incorrect, leading to suboptimal decision-making in practical applications.
In cases where the dataset is particularly small or inadequate, machine learning models—especially complex ones like deep learning architectures—fail to learn the intricate relationships within the data. Complex models require large amounts of data to effectively capture subtle patterns (Keshari et al. 2020). Without sufficient data, the model may not converge, or it may underfit, meaning it performs poorly on both the training and test sets due to an inability to learn the underlying patterns.
@@
Moreover, the use of inadequate data increases variability in model performance. Small datasets are more sensitive to changes in initialization values, training procedures, or even small differences in the data samples used during training (Dodge et al. 2020). As a result, models trained on inadequate data exhibit inconsistent and unpredictable behavior, making them unreliable in different testing or production environments. Issues such as overfitting, bias toward dominant classes, poor handling of noisy data, and a lack of generalizability can result from insufficient, non-representative, or poor-quality data. Ensuring access to high-quality, diverse, and sufficiently large datasets is essential for developing models that are both robust and reliable in their predictions.
@@
Having a reliable means of determining the quantity of training data required to effectively train a model would be an incredibly useful tool for researchers to have. This owes both to the upfront challenges and associated costs of data collection, and with the advent of increasingly costly training runs, the investment costs associated with training.
This is a widely recognized need across research domains, with many experimental protocols requiring power analyses, or equivalent statistical studies to prospectively determine the requisite amount of data collection required to run studies effectively.(Bausell and Li 2002)
However, in spite of the obvious utility of such an assessment, it remains elusive for machine learning model developers. As already noted, many methods for external dataset-based validation but are post-hoc model dependent.
Background
Prospective Analysis of Data Sufficiency
In most academic disciplines, conducting a sufficiently robust study design is arguably more important than the statistical analysis that succeeds it. After-all, a poorly designed study may be unsalvageable after the fact, whereas a poorly analyzed study can simply be re-analyzed once errors in methodology are determined (BMJ 2020).
A critical component in study design is the determination of the appropriate sample size. The sample size must be large enough to establish statistical significance of any potential findings, yet not so large as to unnecessarily burden researchers with the (often costly) acquisition of data.
Attempting to determine data sufficiency for machine learning (ML) models is a particularly vexing challenge given the nature of machine learning. Unlike statistics, ML does not attempt to asset any factual relationship between label and feature set (Bennett et al. 2022); Rather ML models are almost mercenary in nature. Charged with best accomplishing a given prediction or classification task, irrespective of any actual underlying statistical relationship.
This relaxed focus allows for a more expansive inclusion of features for whom a statistically significant relationship to the label may not exist.
Current approaches to determining sufficient sample size
There is a lack of approaches that exist for determining a sufficient sample size to train ML models. The absence of a robust pre-hoc approach that is broadly applicable remains a significant challenge in machine learning, largely stemming from the limited research conducted in this area (Balki et al. 2019).
The most common current approach is empirical analysis. This post-hoc method occurs when the sample size for training is incrementally increased and the learning curve is analyzed to see where the point of diminishing returns occurs in relation to sample size. This enables the identification of the correlation between sample size and model performance.
@@
The only existing approaches that can be considered a pre-hoc approach are model-based. These approaches utilize the parameters of an algorithm to determine sample size. Some models, like neural networks, typically need more data to find a pattern and develop a function that maximizes accuracy when compared to simpler models like decision trees. This is due to the nature of how the model is developed to predict. The more parameters a model has, the more complex it tends to be. Neural networks typically have multiple parameters in each of the hidden layers whereas other algorithms only have one layer. There are mathematical equations like those of Baum and Hausler that allow one to determine a sufficient number of samples through this model-based framework. However, this framework is not optimal for all tasks and has its trade-offs.(Balki et al. 2019)
Statistical Effect size
@@
Effect size is a quantitative measure of the magnitude of a phenomenon. It provides an indication of the practical significance of a result, independent of sample size, making it a critical complement to p-values in hypothesis testing. While p-values indicate whether a result is statistically significant, effect size reveals how strong the effect is, giving a clearer sense of its importance in real-world terms (Sullivan and Feinn 2012).
In the context of a categorical classification modeling challenge, effect size can serve as a critical pre-training tools to compare differences in feature space between classes.
@@
Cohen’s d is commonly used for continuous feature variables, while odds ratios (ORs) are suitable for categorical variables. Both measures help quantify the magnitude of differences across populations.
The formula is
$d=\frac{\mu_{1}-\mu_{2}}{SD_{\text{pooled}}}$ (1)
where
  • $\mu_{1}$ and $\mu_{2}$ are the means of the two populations.
  • The pooled standard deviation ($SD_{\text{pooled}}$) is calculated as
@@
Odds are calculated as the ratio of the probability of an event occurring to the probability of it not occurring:
$\text{Odds}=\frac{p}{1-p}$ (4)
where $p$ is the probability of the event.
This allows us to calculate the individual effect sizes for each feature against a class (Hae-Young 2015). The cumulative effect size of an entire dataset against a class, can then be averaged for a resulting cumulative effect size score.
Hypothesis
@@
Hypothesis 2: There is a correlation between statistical effect-size, data size and model performance up to a point
@@
Methods
For this study, a dataset containing a sampling of adult census data was employed. The dataset was sourced from the UCI Machine Learning Repository and includes 48,842 observations across 14 features.
The following were the features along with data types:
Table 1: Features in the Census dataset.
@@ Methods
Opting for a binary classification problem, the label for this study was income, which was segmented into a binary categorical segmented as either greater than 50k or less than or equal to 50k per year (Becker and Kohavi 1996).
@@
Experiment 1.1.1: Model performance was compared across the same set of features/different data.
For this experiment, the ”Income Greater Than 50k” binary categorical feature served as label, with the rest of the features serving as the feature set. For each data subset, the average effect size was calculated for the binary label we will aim to classify.
The effect size for numerical features was calculated using Cohen’s d, which measures the standardized difference between the means of two groups (e.g., individuals with income greater than \$50K and those with less than or equal to \$50K). For categorical features, the effect size is calculated using the Odds Ratio derived from a contingency table that compares the frequency distribution of the categories across the target variable (income).
After calculating the effect size for each individual feature, the average effect size across all features within a subset is computed. This average provides a single summary metric representing the overall effect size of the dataset in that particular subset.
An identical sequence of classifiers was trained on the models, employing different classification techniques. This was done to try and generalize the approach across learning model paradigms. Specifically, for every experiment, we employed the same four machine-learning models:
  1. 1.
    Logistic Regression
  2. 2.
    Decision Tree
  3. 3.
    Random Forest
  4. 4.
    Neural Network.
This yielded 66 data points of model performance vs. averaged effect size, allowing us to compare those two elements by calculating a direct correlation.
@@
Experiment 1.2.1: For this experiment, model performance across a differing set of features with the same underlying data was compared. In this experiment the ”income greater than 50k” binary was retained as a label, while the rest of the features were retained as a feature set. For this experiment, a series of subsets were devised. For each subset one of the features was dropped, and then the averaged effect size for the binary label was recalculated on the reduced feature space. An identical sequence of classifiers was then trained, and model performance and associated averaged effect size were calculated.
This produced 14 data points of model performance vs. effect size.
@@
This experiment sought to determine if there a relationship between the slope of a learning curve (the rate at which a model’s error rate decreases with respect to training data size) and the effect size, irrespective of overall model performance. The hypothesis being tested was whether or not a more dramatic effect size would indicate a cleaner dataset, requiring fewer samples for a model to ‘bottom out’ in terms of error. The critical distinction here was that it would not necessarily predict model performance, but rather how large a dataset was required for convergence on an optimal model.
For each model in experiment 1, where a previously calculated effect size was obtained, a learning curve/error rate for both training and validation sets was plotted.
@@
Results
Figure 1: Experiment 1.1.1 (income as label) with $R^{2}=0.0132$.
Figure 2: Experiment 1.1.2 (gender as label) with $R^{2}=0.0011$.
Figure 3: Experiment 1.2.1 (income as label) with $R^{2}=0.0142$.
Figure 4: Experiment 1.2.2 (gender as label) with $R^{2}=0.091$.
Experiment 1
Table 2: Summary of results for Experiment 1.
Exp. No. Experiment Name R2 Value
1.1.1 Income label, identical features, different subsets 0.0132
1.1.2 Sex label, identical features, different subsets 0.0011
1.2.1 Income label, different feature mix 0.0142
1.2.2 Sex label, different feature mix 0.091
Table 2 details the experimental results.
Overall the results demonstrably show that effect size has essentially no predictive value on the resulting model’s performance. As demonstrated by the corresponding R2 values, virtually none of the variance in model performance is attributable to statistical effect size.
While this work does not preclude the possibility of alternative statistical measures possibly serving as a metric for resulting model performance, thus far, such pre-hoc assessments remain elusive.
As ML algorithms are designed to maximize the performance of their outputs and minimize error irrespective of the input data, effective learning mechanisms can employ nonlinear and nonstatistical approaches to modeling data that cannot be effectively determined pre-hoc using relatively straight-forward linear statistical measures. Ultimately, this experiment suggests that model performance may have to be determined by some other factor of the dataset along with the specific model utilized.
Figure 5: Experiment 2.1 with $R^{2}=0.0054$.
Figure 6: Experiment 2.2 with $R^{2}=0.0007$.

### v3

@@ Abstract
  1. 1.
  2. 2.
-A Introduction
@@
The sufficiency of a dataset, both in terms of size and representativeness, is critical to training an effective Machine Learning model. Failure to train on data of adequate quality is likely to result in models that dramatically under-perform in production environments [7].
@@
Training machine learning models on inadequate data presents several challenges that can significantly impact model performance and generalizability. While the problem of ’inadequacy’ of data is a multifaceted one, one significant component of it is a lack of sufficient data volume, which in-turn likely impacts the representativeness of the data sample towards the overall population undermining the ability of a model to learn meaningful patterns, resulting in overfitting, where the model performs well on the training data but struggles to generalize to unseen data leading to poor outcomes when deployed in real-world scenarios. This owes to the fact that if the dataset used for training does not accurately reflect the conditions and distributions the model will encounter in real-world applications, the model will struggle to generalize beyond the narrow scope of its training [10]. Consequently, the model’s assumptions about the domain may be incorrect, leading to suboptimal decision-making in practical applications.
@@
In cases where the dataset is particularly small or inadequate, machine learning models—especially complex ones like deep learning architectures—fail to learn the intricate relationships within the data. Complex models require large amounts of data to effectively capture subtle patterns [9]. Without sufficient data, the model may not converge, or it may underfit, meaning it performs poorly on both the training and test sets due to an inability to learn the underlying patterns. Moreover, the use of inadequate data increases variability in model performance. Small datasets are more sensitive to changes in initialization values, training procedures, or even small differences in the data samples used during training [6]. As a result, models trained on inadequate data exhibit inconsistent and unpredictable behavior, making them unreliable in different testing or production environments. Issues such as overfitting, bias toward dominant classes, poor handling of noisy data, and a lack of generalizability can result from insufficient, non-representative, or poor-quality data. Ensuring access to high-quality, diverse, and sufficiently large datasets is essential for developing models that are both robust and reliable in their predictions.
@@
Having a reliable means of determining the quantity of training data required to effectively train a model would be an incredibly useful tool for researchers to have. This owes both to the upfront challenges and associated costs of data collection, and with the advent of increasingly costly training runs, the investment costs associated with training. This is a widely recognized need across research domains, with many experimental protocols requiring power analyses, or equivalent statistical studies to prospectively determine the requisite amount of data collection required to run studies effectively.[2] However, in spite of the obvious utility of such an assessment, it remains elusive for machine learning model developers. As already noted, many methods for external dataset-based validation but are post-hoc model dependent.
-B Background
-B1 Prospective Analysis of Data Sufficiency
In most academic disciplines, conducting a sufficiently robust study design is arguably more important than the statistical analysis that succeeds it. After-all, a poorly designed study may be unsalvageable after the fact, whereas a poorly analyzed study can simply be re-analyzed once errors in methodology are determined [5]. A critical component in study design is the determination of the appropriate sample size. The sample size must be large enough to establish statistical significance of any potential findings, yet not so large as to unnecessarily burden researchers with the (often costly) acquisition of data. Attempting to determine data sufficiency for machine learning (ML) models is a particularly vexing challenge given the nature of machine learning. Unlike statistics, ML does not attempt to asset any factual relationship between label and feature set [4]; Rather ML models are almost mercenary in nature. Charged with best accomplishing a given prediction or classification task, irrespective of any actual underlying statistical relationship. This relaxed focus allows for a more expansive inclusion of features for whom a statistically significant relationship to the label may not exist.
-B2 Current approaches to determining sufficient sample size
@@
There is a lack of approaches that exist for determining a sufficient sample size to train ML models. The absence of a robust pre-hoc approach that is broadly applicable remains a significant challenge in machine learning, largely stemming from the limited research conducted in this area [1]. The most common current approach is empirical analysis. This post-hoc method occurs when the sample size for training is incrementally increased and the learning curve is analyzed to see where the point of diminishing returns occurs in relation to sample size. This enables the identification of the correlation between sample size and model performance. The only existing approaches that can be considered a pre-hoc approach are model-based. These approaches utilize the parameters of an algorithm to determine sample size. Some models, like neural networks, typically need more data to find a pattern and develop a function that maximizes accuracy when compared to simpler models like decision trees. This is due to the nature of how the model is developed to predict. The more parameters a model has, the more complex it tends to be. Neural networks typically have multiple parameters in each of the hidden layers whereas other algorithms only have one layer. There are mathematical equations like those of Baum and Hausler that allow one to determine a sufficient number of samples through this model-based framework. However, this framework is not optimal for all tasks and has its trade-offs.[1]
-B3 Statistical Effect size
@@
Effect size is a quantitative measure of the magnitude of a phenomenon. It provides an indication of the practical significance of a result, independent of sample size, making it a critical complement to p-values in hypothesis testing. While p-values indicate whether a result is statistically significant, effect size reveals how strong the effect is, giving a clearer sense of its importance in real-world terms [11]. In the context of a categorical classification modeling challenge, effect size can serve as a critical pre-training tools to compare differences in feature space between classes.
@@
Cohen’s d is commonly used for continuous feature variables, while odds ratios (ORs) are suitable for categorical variables. Both measures help quantify the magnitude of differences across populations. The formula is $d=\frac{\mu_{1}-\mu_{2}}{SD_{\text{pooled}}},$ where $\mu_{1}$ and $\mu_{2}$ are the means of the two populations, and the pooled standard deviation ($SD_{\text{pooled}}$) is calculated as
@@
Odds are calculated as the ratio of the probability of an event occurring to the probability of it not occurring $\text{Odds}=\frac{p}{1-p},$ where $p$ is the probability of the event. This allows us to calculate the individual effect sizes for each feature against a class [8]. The cumulative effect size of an entire dataset against a class, can then be averaged for a resulting cumulative effect size score.
-C Hypothesis
@@
Figure 1: Experiment 1.1.1 (income as label): Mean absolute effect size plotted against model performance (accuracy, precision, recall, and F1-score). Each point represents a model (distinguished by color) trained on a distinct subset.
@@
Hypothesis 2: There is a correlation between statistical effect-size, data size, and model performance up to a point.
@@
I Methods
For this study, a dataset containing a sampling of adult census data was employed. The dataset was sourced from the UCI Machine Learning Repository and includes 48,842 observations across 14 features. The features along with data types are shown at Table I.
TABLE I: Features in the Census dataset.
@@ I Methods
Figure 2: Experiment 1.1.2 (gender as label): Mean absolute effect size plotted against model performance (accuracy, precision, recall, and F1-score). Each point represents a model (distinguished by color) trained on a distinct subset.
Figure 3: Experiment 1.2.1 (income as label): Effect size of removed feature plotted against model performance drop (accuracy drop, precision drop, recall drop, and F1-score drop). Each point represents a model (distinguished by color) with a different feature removed.
Figure 4: Experiment 1.2.2 (gender as label): Effect size of removed feature plotted against model performance drop (accuracy drop, precision drop, recall drop, and F1-score drop). Each point represents a model (distinguished by color) with a different feature removed.
Figure 5: Experiment 2.1 (left): Mean absolute effect size plotted against slope of the learning curve. Experiment 2.2 (right): Mean absolute effect size plotted against slope of the difference between training learning curve and validation learning curve. Both experiments set income as label. Each point represents a model (distinguished by color) trained on a distinct subset.
Opting for a binary classification problem, the label for this study was income, which was segmented into a binary categorical segmented as either greater than 50k or less than or equal to 50k per year [3].
@@ I-1 Experiment 1
Experiment 1.1.1: Model performance was compared across the same set of features/different data. For this experiment, the ”Income Greater Than 50k” binary categorical feature served as label, with the rest of the features serving as the feature set. For each data subset, the average effect size was calculated for the binary label we will aim to classify. The effect size for numerical features was calculated using Cohen’s d, which measures the standardized difference between the means of two groups (e.g., individuals with income greater than \$50K and those with less than or equal to \$50K). For categorical features, the effect size is calculated using the Odds Ratio derived from a contingency table that compares the frequency distribution of the categories across the target variable (income).
After calculating the effect size for each individual feature, the average effect size across all features within a subset is computed. This average provides a single summary metric representing the overall effect size of the dataset in that particular subset. An identical sequence of classifiers was trained on the models, employing different classification techniques. This was done to try and generalize the approach across learning model paradigms. Specifically, for every experiment, we employed the same four machine-learning models: Logistic Regression, Decision Tree, Random Forest, and Neural Network. This yielded 66 data points of model performance vs. averaged effect size, allowing us to compare those two elements by calculating a direct correlation.
@@
Experiment 1.2.1: For this experiment, model performance across a differing set of features with the same underlying data was compared. In this experiment the ”income greater than 50k” binary was retained as a label, while the rest of the features were retained as a feature set. A baseline on all performance metrics was initially calculated. For this experiment, a series of subsets were devised. For each subset one of the features was dropped, and the effect size of the dropped feature was calculated. An identical sequence of classifiers was then trained, and model performance drop was calculated. This produced 14 data points of model performance drop vs. effect size of removed feature.
@@ I-2 Experiment 2
This experiment sought to determine if there a relationship between the slope of a learning curve (the rate at which a model’s error rate decreases with respect to training data size) and the effect size, irrespective of overall model performance. The hypothesis being tested was whether or not a more dramatic effect size would indicate a cleaner dataset, requiring fewer samples for a model to ‘bottom out’ in terms of error. The critical distinction here was that it would not necessarily predict model performance, but rather how large a dataset was required for convergence on an optimal model. For each model in experiment 1, where a previously calculated effect size was obtained, a learning curve/error rate for both training and validation sets was plotted.
@@
II Results
Multiple inferential metrics were utilized to evaluate results. Pearson correlation (r with associated p-value) to evaluate linear relationship. The r-squared value was derived to determine how much of the variance in the dependent variable can be explained by the independent variable. Spearman correlation ($\rho$ with associated p-value) to detect monotonic relationship. Kendall’s Tau ($\tau$ with associated p-value) to establish whether any consistent ordering between variables exists. A 95% bootstrap confidence interval for pearson and spearman was also implemented for further validation of results, since experiments were run on a single dataset. Here the data was resampled 1000 times to see if any correlation that does exist is sound based on if it shows up consistently. Table II details the experimental results and plots for each experiment can be found in Figures 1,2, 3, 4, and 5.
TABLE II: Complete Summary of Inferential Metrics Across All Experiments
Exp. Metric Pearson $r$ $p$ $R^{2}$ Spearman $\rho$ $p$ Kendall $\tau$ $p$ Slope Pearson CI Spearman CI
Experiment 1.1.1 — Income Label, Identical Features
1.1.1 Accuracy -0.083 0.181 0.007 -0.052 0.399 -0.035 0.417 -0.028 [-0.205, 0.047] [-0.172, 0.067]
1.1.1 Precision -0.056 0.362 0.003 -0.033 0.590 -0.024 0.568 -0.059 [-0.164, 0.054] [-0.152, 0.088]
1.1.1 Recall 0.101 0.102 0.010 0.121 0.049 0.080 0.057 0.101 [-0.013, 0.217] [0.007, 0.241]
1.1.1 F1 Score 0.061 0.327 0.004 0.078 0.204 0.049 0.236 0.051 [-0.058, 0.171] [-0.038, 0.198]
Experiment 1.1.2 — Sex Label, Identical Features
1.1.2 Accuracy -0.173 0.005 0.030 -0.150 0.014 -0.101 0.018 -0.072 [-0.287, -0.056] [-0.272, -0.038]
1.1.2 Precision -0.139 0.024 0.019 -0.123 0.046 -0.084 0.045 -0.051 [-0.254, -0.014] [-0.238, 0.001]
1.1.2 Recall -0.145 0.018 0.021 -0.086 0.165 -0.057 0.169 -0.065 [-0.265, -0.027] [-0.204, 0.025]
1.1.2 F1 Score -0.185 0.003 0.034 -0.144 0.019 -0.097 0.021 -0.058 [-0.305, -0.066] [-0.263, -0.029]
Experiment 1.2.1 — Income Label, Feature Removal
1.2.1 Accuracy Drop -0.123 0.365 0.015 -0.003 0.982 -0.017 0.864 -0.001 [-0.387, 0.164] [-0.263, 0.275]
1.2.1 Precision Drop -0.138 0.312 0.019 0.012 0.930 -0.009 0.926 -0.004 [-0.402, 0.157] [-0.254, 0.285]
1.2.1 Recall Drop -0.118 0.387 0.014 -0.055 0.688 -0.043 0.666 -0.004 [-0.373, 0.111] [-0.326, 0.232]
1.2.1 F1 Drop -0.141 0.302 0.020 -0.039 0.774 -0.031 0.744 -0.004 [-0.406, 0.143] [-0.330, 0.249]
Experiment 1.2.2 — Sex Label, Feature Removal
1.2.2 Accuracy Drop -0.068 0.621 0.005 -0.154 0.258 -0.105 0.286 -0.001 [-0.344, 0.202] [-0.429, 0.132]
1.2.2 Precision Drop -0.026 0.850 0.001 -0.141 0.301 -0.093 0.328 -0.000 [-0.207, 0.204] [-0.397, 0.135]
1.2.2 Recall Drop -0.053 0.700 0.003 -0.055 0.690 -0.038 0.702 -0.000 [-0.289, 0.192] [-0.318, 0.221]
1.2.2 F1 Drop -0.071 0.606 0.005 -0.150 0.271 -0.097 0.307 -0.000 [-0.368, 0.220] [-0.415, 0.142]
Experiment 2.1 — Learning Curve Slope
2.1 LC Slope -0.115 0.099 0.013 -0.185 0.007 -0.129 0.006 -0.042 [-0.274, 0.049] [-0.326, -0.031]
Experiment 2.2 — Train–Validation Gap Slope
2.2 Gap Slope 0.065 0.349 0.004 0.103 0.138 0.072 0.124 0.000 [-0.097, 0.206] [-0.049, 0.236]
II-A Experiment 1
Experiment 1.1.1: The pearson r-values do not indicate any relationship and the corresponding p-values indicate that any relationship that may exist is statistically insignificant. Figure 1 is consistent with this result and shows no linear trend in the data. The spearman-rho value for recall suggests a very weak positive monotonic relationship. The corresponding p-value indicates that although the relationship is weak, it is likely not a random result. The spearman for the other performance metrics does not reveal any statistically detectable relationships. The slope for recall is approximately 10% and the spearman confidence interval for recall excludes zero. These results align with the positive monotic relationship observed and confirm consistency.
Experiment 1.1.2: The pearson r-values and corresponding p-values indicate that a linear negative relationship is statistically detectable, but is very weak. As the absolute average effect size increases the performance seems to slightly decrease. We see in Accuracy and F1-Score this relationship is slightly stronger. The spearman-$\rho$ values for Accuracy, Precision, and F1-Score suggest a very weak negative monotonic relationship. And the corresponding p-values for these indicate that although the relationship is weak, it is likely not a random result. The kendall-tau and slope values are quite small, but they do align with the negative trend observed. In the confidence intervals, the pearson excludes zero for all performance metrics. Further confirming that although the relationship is weak, it is consistent. Similarly with Spearman confidence intervals, zero is excluded for Accuracy and F1-Score. Figure 2 visually reflects these results as there is a very slight downtrend when compared to to Figure 1.
Experiment 1.2.1: The pearson r-values suggest a weak negative correlation, but corresponding p-values are very high. This indicates that statistically we cannot reject the null hypothesis. Figure 3 and the remaining metrics do not suggest any statistically detectable relationships between the effect size of the removed feature and the change in performance.
Experiment 1.2.2 The pearson r-values and kendall values show no relationship. The spearman $\rho$-values for accuracy drop, precision drop, and F1 drop signal a weak negative monotonic relationship. However, the corresponding p-values for these three inferential metrics confirm that they are statistically insignificant. Figure 4 and the confidence intervals align with these results, showing no detectable consistent relationship.
II-B Experiment 2
Experiment 2.1: The pearson r-value for the experiment indicates a negative correlation, but the p-value considers this statistically insignficant. The spearman $\rho$-value suggests a weak negative montonic relationship between absolute average effect size and the slope of the learning curve. The corresponding p-value confirms this relationship is unlikely due to random chance. The kendall value and its p-value support this observation and suggest a very weak relationship as well. As absolute average effect size goes up, the slope of the learning curve slightly decreases. The spearman confidence interval confirms consistency.